# DrugMCTS: a drug repurposing framework combining multi-agent, RAG and Monte Carlo Tree Search

### v2

@@ 5 Conclusion
Our DrugMCTS framework revolutionizes drug repurposing by integrating multi-agent collaboration (five specialized agents), hybrid data processing (scientific → hybrid → general), and Monte Carlo Tree Search to enable the lightweight Qwen2.5-7B-Instruct model to outperform Deepseek-R1 by more than 20% recall on DrugBank/KIBA datasets. The system achieves 55.34% recall via dynamic Top-K+3 selection, validated by 1,221 experimental interactions and case studies like Equol-CXCR3 binding (docking score: -8.4 kcal/mol). This work establishes a template for LLM-powered scientific discovery beyond drug-target prediction.
@@ 6 Limitations
Appendix A Appendix
A.1 Dataset Construction
@@
Table 3: Dataset Statistics Summary
@@
Table 5: Class Distribution Statistics

### v3

@@ 4.1 Settings
Since excluding certain actions in MCTS can occasionally reduce the number of generated nodes, sometimes below the number of ground-truth targets, we ensure fair comparison by setting the MCTS-based baselines to generate four candidate nodes during the Protein Selection Action, instead of a single node used in the full pipeline.
@@ 5 Conclusion
Our DrugMCTS framework revolutionizes drug repositioning by integrating multi-agent collaboration (five specialized agents), hybrid data processing (scientific → hybrid → general), and MCTS to enable the lightweight Qwen2.5-7B-Instruct model to outperform Deepseek-R1 by more than 20% recall on DrugBank/KIBA datasets. The system achieves 55.34% recall via dynamic Top-K+3 selection, validated by 1,221 experimental interactions and case studies like Equol-CXCR3 binding (docking score: -8.4 kcal/mol). This work establishes a template for LLM-powered scientific discovery beyond drug-target prediction.
@@ 6 Limitations
Appendix 0.A Appendix
Algorithm 1 A single MCTS rollout within the DrugMCTS framework.
1:Query molecule $M_{qm}$
2:Selected protein $P_{s}$, final reward $R_{\text{final}}(p^{*})$
3:// A1 Retrieval Action
4:Compute Tanimoto and ChemBERTa cosine similarities
[…]
7:// A2 Molecule Analysis Action
[…]
10:// A3 Molecule Selection Action
11:For all $m\in M_{cm}$: extract $(C_{c,s},C_{c,phy})$ using MA agent
[…]
14:// A4 Interaction Analysis Action
[…]
18:// A5 Protein Selection Action
[…]
20:// A6 End Action
21:Perform backpropagation in MCTS from end node
22:// Reward Calculation
23:Query model $k$ times to get selections $\{p^{(1)},\dots,p^{(k)}\}$
[…]
26:Feed $(P_{s},D_{bp},L_{rp},M_{qm},R_{qm})$ to decision model $k$ times
[…]
28:Return: $P_{s}$, $R_{\text{final}}(p^{*})=\frac{R_{\text{relative}}(p^{*})+R_{\text{absolute}}(p^{*})}{2}$
Algorithm 2 MCTS for DrugMCTS
1:Query molecule $M_{qm}$, number of simulations $N$
2:Selected protein $P_{s}$
3:Initialize root node $s_{0}$ with $M_{qm}$
4:for $i=1$ to $N$ do
[…]
8:  Backpropagate reward $R_{i}$ to update tree
9:end for
10:return $P_{s}\leftarrow\textbf{BestChild}(s_{0})$ based on average reward
0.A.1 Dataset Construction
@@
Table 3: Experimental Dataset Statistics
@@
Table 5: Deep Learning Dataset Statistics